OpenStudy (anonymous):

Integrate (e^-x - e^x)/(e^2x + e^-2x +2) dx

OpenStudy (zzr0ck3r):

start with u = e^x then you will have partial fractions to deal with, then more substitutions on each term

OpenStudy (anonymous):

let u=e^x+e^(-x)

OpenStudy (anonymous):

\[u=e^x+e^{-x} \ \ \ then \ \ \ du=(e^x-e^{-x})dx \\ e^{2x}+e^{-2x}+2=(e^x+e^{-x})^2=u^2 \\ so \ \ I=\int\limits \frac{-du}{u^2}\]

OpenStudy (anonymous):

notice e^-2x+e^-2x+2=(e^x+e^-x)^2 also (e^-x-e^x)/(e^x+e^-x)^2=(1/2)((e^-x-e^x)/2/((e^x+e^-x)^2)/4) =(1/2)(-(e^x-e^-x)/2/(coshx)^2)=(1/2)(-sinh x/(cosh x)^2) so we have: \[-1/2\int\limits sinhx/(coshx)^2dx\] use sub. u=coshx du=sinhxdx so \[-1/2\int\limits 1/u^2du=1/2u+c=1/2\cosh(x) + c\]

OpenStudy (anonymous):

\[I=\int\limits \frac{-du}{u^2}=\frac{1}{u}+c=\frac{1}{e^x+e^{-x}}+c=\frac{1}{2 \cosh x}+c\]

OpenStudy (anonymous):

I really feel like if he needs help with this hyperbolic trig functions aren't going to help...most people have trouble enough with ordinary trig functions lol...